# Deep Learning Approach for Changepoint Detection: Penalty Parameter Optimization

### v2

@@ 4 $\leq 64$ 1 1 2 2 1 1 1 9
For each model MLP from each dataset, 6 folds are tested, resulting in the selection of 6 MLP configurations. Dataset detailed has 3,730 sequences, systematic has 3,418, and epigenomic has 4,913. Observing the trend in the table, it’s evident that the preference leans towards MLPs with 2 or 3 hidden layers and the number of neurons is less than 64. Specifically, out of the 72 chosen architectures, 15 have 1 hidden layer, 25 have 2 hidden layers, 19 have 3 hidden layers, and 13 have 4 hidden layers.
@@ Discussion:
Effect of Feature Selection: The experiments reveal that accuracy improves when a larger set of features (except in the case of 117 features) is integrated, compared to using fewer features, across linear, decision tree, or MLP models. Perhaps the reason lies in the better quality of the larger feature set. If we closely examine the feature sequence length in Figure 5, the relationship between sequence length and the target interval of $\lambda$ is not entirely clear. This lack of clarity could explain why models relying solely on this feature exhibit lower accuracy compared to others. Remarkably, nonlinear models such as decision trees or MLPs exhibit lower accuracy when employing 117 features compared to when using only 4 features. So selecting only 4 features proves highly beneficial, enhancing model accuracy (both linear models and MLPs) while also potentially reducing training time for the models. This divergence can be attributed to the inherent challenge faced by MLPs in effectively eliminating unnecessary features.
@@
Using MLPs effect: employing MLPs with appropriate configuration (see Figure 7 and Table 2) yields a enhancement in accuracy when contrasted with linear models. In the case of the same feature set, it’s not guaranteed that MLPs will consistently outperform linear models. For instance, in the dataset detailed, the linear model with two features perform better than the MLP. Similarly, in the dataset systematic, the linear model with one feature has the better performance than the MLP.
Comparison of MMIT and Linear Model Accuracy: In many cases where the number of features remains constant, MMIT does not outperform linear models. This is often due to observable linear relationships between features and target intervals (see Figure 5). Intuitively, linear models are perceived as superior to decision trees under such circumstances.
@@
Conclusion: MLP models with chosen four features generally achieve higher accuracy in changepoint detection compared to linear models and decision trees. However, it’s important to note that the training process for MLPs is significantly more time-consuming than that for linear models.
@@
Limitations of this study: There are several considerable limitations from this study below
@@
May not be applicable to other sequence datasets: Intuitively, features such as variance, value range, and sum of absolute difference play a role in predicting this penalty problem for these particular benchmark datasets. However, these features may not be relevant in other types of sequence datasets.
@@
May overlook some other useful features: By focusing on a limited number of features and their relationship with target intervals, we may overlook additional useful features and potential interactions between features.
@@
Challenges in Selecting the Optimal MLP Configuration The process of identifying the most suitable MLP configuration is inherently time-consuming. It entails evaluating a plethora of configurations, each comprising different combinations of hidden layers and neurons. Our task entails validating 36 distinct MLP models for each pair of train and test sets (we considered variations in the number of hidden layers, ranging from 1 to 4, and the number of neurons per layer, which can be 2, 4, 8, 16, 32, 64, 128, 256, or 512). This exhaustive exploration represents a significant computational challenge. Furthermore, we did not delve into exploring architectures with varying neuron counts across layers within a single model.
@@
Future Work: Various neural network architectures can be explored. Beyond MLPs, one might investigate the performance of other architectures tailored to the specific characteristics of the data. In some cases, these alternative network structures may yield superior results compared to MLPs or linear models. About feature selection, instead of manual feature extraction, it could be worthwhile to explore Recurrent Neural Networks (RNNs) (Hopfield, 1982) for directly extracting features from raw sequences. Variants like Gated Recurrent Units (GRUs) (Cho et al., 2014) or Long Short-Term Memory networks (LSTMs) (Hochreiter and Schmidhuber, 1997) could be examined for this task. Moreover, as feature transformations like logarithmic scaling were employed, it’s worth exploring alternative approaches to feature engineering.
@@
Reproducible Research Material: For those interested in replicating our study, all the code and associated materials are available at this link:

### v3

@@ 4 $\leq 64$ 1 1 2 2 1 1 1 9
00footnotetext: For each MLP model from each dataset, 6 folds are tested, resulting in the selection of 6 MLP configurations. The dataset ’detailed’ has 3,730 sequences, ’systematic’ has 3,418, and ’epigenomic’ has 4,913. Observing the trend in the table, it is evident that the preference leans towards MLPs with 2 or 3 hidden layers and not greater than 64 neurons. Specifically, out of the 72 chosen architectures, 15 have 1 hidden layer, 25 have 2 hidden layers, 19 have 3 hidden layers, and 13 have 4 hidden layers.
@@ Discussion
The experiments show that accuracy improves as the number of features increases (from 1 to 2 to 4, excluding the case of all features), regardless of whether linear, tree-based, or MLP models are used. Perhaps the reason lies in the better quality of the larger feature set. If we closely examine the feature sequence length in Figure 5, the relationship between sequence length and the target interval of $\lambda$ is not entirely clear. This lack of clarity could explain why models relying solely on this feature exhibit lower accuracy compared to others. Remarkably, nonlinear models such as decision trees or MLPs exhibit lower accuracy when employing all features compared to when using only 4 features. So selecting only 4 features proves highly beneficial, enhancing model accuracy (both linear models and MLPs) while also potentially reducing training time for the models. This divergence can be attributed to the inherent challenge faced by MLPs in effectively eliminating unnecessary features.
Using MLPs effect
@@
employing MLPs with appropriate configuration (see Figure 7 and Table 2) yields a enhancement in accuracy when contrasted with linear models and trees. In the case of the same feature set, it’s not guaranteed that MLPs will consistently outperform linear models. For instance, in the dataset detailed, the linear model with two features perform better than the MLP. Similarly, in the dataset systematic, the linear model with one feature has the better performance than the MLP.
Comparison of MMIT, AFT in XGBoost vs Linear Model Accuracy
In many scenarios where the number of features remains constant, MMIT and AFT in XGBoost do not exhibit superior performance compared to linear models. This observation is consistent with results reported in [14] for the dataset Systematic, and [15] for the dataset Epigenomic, where the accuracy rates of AFT in XGBoost were similar to those of linear models and MMIT. This lack of improvement is often attributed to the presence of observable linear relationships between features and target intervals, as illustrated in Figure 5.
Comparison of MLP and Linear Model Accuracy
Examine Figure 6, for the linear model, performance generally improves as the number of features increases. However, this trend does not hold for MLP models; when there are too many features, their performance declines. Additionally, with just 1 or 2 features, MLP does not outperform the linear model—MLP only surpasses the linear model when 4 features are used.
Conclusion
@@
MLP models with chosen four features generally achieve higher accuracy in changepoint detection compared to linear models and decision trees. However, it’s important to note that the training process for MLPs is significantly more time-consuming than that for linear models.
Limitations of this study
@@
There are several considerable limitations from this study below
May not be applicable to other sequence datasets
@@
Intuitively, features such as variance, value range, and sum of absolute difference play a role in predicting this penalty problem for these particular benchmark datasets. However, these features may not be relevant in other types of sequence datasets.
May overlook some other useful features
@@
By focusing on a limited number of features and their relationship with target intervals, we may overlook additional useful features and potential interactions between features.
Challenges in Selecting the Optimal MLP Configuration
@@
The process of identifying the most suitable MLP configuration is inherently time-consuming. It entails evaluating a plethora of configurations, each comprising different combinations of hidden layers and neurons. Our task entails validating 36 distinct MLP models for each pair of train and test sets (we considered variations in the number of hidden layers, ranging from 1 to 4, and the number of neurons per layer, which can be 2, 4, 8, 16, 32, 64, 128, 256, or 512). This exhaustive exploration represents a significant computational challenge. Furthermore, we did not delve into exploring architectures with varying neuron counts across layers within a single model.
Future Work
@@
Various neural network architectures can be explored. Beyond MLPs, one might investigate the performance of other architectures tailored to the specific characteristics of the data. In some cases, these alternative network structures may yield superior results compared to MLPs or linear models. About feature selection, instead of manual feature extraction, it could be worthwhile to explore Recurrent Neural Networks (RNNs) [29] for directly extracting features from raw sequences. Variants like Gated Recurrent Units (GRUs) [30] or Long Short-Term Memory networks (LSTMs) [31] could be examined for this task. Moreover, as feature transformations like logarithmic scaling were employed, it’s worth exploring alternative approaches to feature engineering.
Reproducible Research Material
@@
For those interested in replicating our study, all the code and associated materials are available at this link:
@@
Declarations
No funding was received for conducting this study.